# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 2.2. Assumptions
((AC))ν $P\circ\mathsf{X}_{t}^{-1}$ is absolutely continuous with respect to $\nu$ for all $0<t<\infty$.
Intuitively, these assumptions mean the following. Condition ((SIN)) ensures that the particle system almost surely remains in the well-behaved state space $W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})$. Condition ((NBJ)) guarantees that within any finite time interval $[0,T]$, only finitely many particles enter any given finite region $R,R]^{d}$, thus preventing an infinite influx of particles or jumps. Condition ((AC))ν ensures that the distribution of particles at each time $t$ admits a density with respect to the reference measure $\nu$, making the distribution comparable to $\nu$.
In addition, we require two further conditions, ((MF)) and ((IFC)), which are described in Subsection 13.5 and Subsection 13.4, respectively. For the precise definitions of a strong solution, a unique strong solution, and pathwise uniqueness of strong solutions, see Definitions 13.1113.13.
@@
It is well known that the pathwise uniqueness of strong solutions ensures the uniqueness in law of weak solutions (see [11]). For the concepts of uniqueness in law and pathwise uniqueness of weak solutions, see Definition 13.8 and Definition 13.9. Based on Theorem 13.5, the following corollary of Theorem 2.3 is obtained.
Corollary 2.
@@
From (2.36) and Theorem 2.3, we immediately obtain the following.
Corollary 3.
Assume ((A1))((A4)). Then, the results in Theorem 2.3 apply to (2.49). The solutions of both (2.41) and (2.49) are indistinguishable if they have the same initial starting point.
Remark 2.8.
Note that $\mu=\int\mu_{\mathsf{a}}\mu(d\mathsf{a})$. A strong solution is a weak solution by definition. Hence, all solutions starting at $\mathfrak{l}(\mathsf{s})$ in Theorem 2.3 give weak solutions for $\mu$-a.s. $\mathsf{s}$.
@@
The analysis of fully labeled dynamics plays a crucial role in understanding the structure of infinite-particle systems. While the $m$-labeled dynamics $\mathbf{X}^{[m]}$ provide a useful finite-dimensional projection that admits an invariant distribution, they capture only a partial view of the system. To obtain a complete picture, one must go beyond these finite labels and establish the existence of a genuine diffusion on the full label space $(\mathbb{R}^{d})^{\mathbb{N}}$. This step is indispensable, since it ensures that the dynamics are well-defined not merely in restricted subsystems but for the entire infinite system.
Historically, the construction of fully labeled dynamics has been regarded as a highly challenging problem. Even in the context of determinantal processes such as the Dyson model, establishing such dynamics required substantial effort, with delicate arguments based on random matrix theory. For general Coulomb systems, the situation is even more severe: the long-range nature of the interaction and the absence of classical Gibbsian structures prevent the direct application of Dirichlet form methods or invariant measure techniques. Against this backdrop, proving the existence of fully labeled conservative diffusions represents a significant advance, as it overcomes obstacles that have persisted for decades.
In Theorem 2.4, we prove that the fully labeled process $\mathbf{X}$ is indeed an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued conservative diffusion. Unlike $\mathbf{X}^{[m]}$, however, the process $\mathbf{X}$ possesses no invariant measure, reflecting a fundamental difference between finite-label and full-label dynamics. Nevertheless, the result guarantees that the fully labeled dynamics form a legitimate stochastic process with infinite lifetime, thereby providing a rigorous foundation for the dynamical theory of Coulomb systems.
As before, for $\mathbf{s}=(s^{i})_{i}$ we write $\mathfrak{u}(\mathbf{s})=\sum_{i}\delta_{s^{i}}$. Let $\mathbf{S}_{\star\star}\subset(\mathbb{R}^{d})^{\mathbb{N}}$ denote the subset defined in (11.13).
@@
Assume ((A1))((A3)). Let $\mathbf{X}$ be the solution of the ISDE (2.41) constructed in Theorem 2.3. Then there exists a measurable subset $\mathbf{S}_{\star\star}\subset(\mathbb{R}^{d})^{\mathbb{N}}$ such that $\mathbf{X}$ is a conservative diffusion with state space $\mathbf{S}_{\star\star}$ and $\mu(\mathfrak{u}(\mathbf{S}_{\star\star}))=1$.
@@
$P_{\mathbf{x}}:=P\big{(}\mathbf{X}\in\cdot\,\big{|}\,\mathbf{X}_{0}=\mathbf{x}\big{)}$
@@
admits a version such that $\{P_{\mathbf{x}}\}_{\mathbf{x}\in\mathbf{S}_{\star\star}}$ forms a family of probability measures on $C([0,\infty);\mathbf{S}_{\star\star})$ with the strong Markov property and infinite lifetime [5]. In particular, the diffusion law $P_{\mathbf{x}}$ is defined for every $\mathbf{x}\in\mathbf{S}_{\star\star}$.
Remark 2.9.
@@
i The state space can in principle be extended to the whole $(\mathbb{R}^{d})^{\mathbb{N}}$ by freezing the dynamics outside $\mathbf{S}_{\star\star}$. The set $\mathbf{S}_{\star\star}$ is given explicitly by (11.13).
ii It is plausible that the diffusion admits no invariant probability measure. An analogous statement is known for the Dyson model [36]. In particular, this suggests that there may be no Dirichlet form naturally associated with the diffusion.
2.7. Finite-volume approximation: Theorem 2.5
Next, we consider a finite-particle approximation of dynamics. We impose reflecting boundary conditions on $\partial S_{R}$ and study the dynamics of the particle systems $\underline{\mathbf{X}}_{R}^{N}$ and $\underline{\mathbf{X}}_{R}$, in which particles outside the region $\overline{S}_{R}$ are held fixed.
Let $\mathsf{d}^{N}$ be the logarithmic derivative of $\mu^{N}$. We write $\mathsf{s}=\sum_{i=1}^{N-1}\delta_{s^{i}}$. From (2.18),
$\mathsf{d}^{N}({x},\mathsf{s})=-\beta\Big{(}\nabla\Phi^{N}({x})+\sum_{i=1}^{N-1}\nabla_{{x}}\Psi^{N}({x},s^{i})\Big{)}.$
@@
Let $\mathfrak{b}^{N}=\{\nabla\mathfrak{a}+\mathfrak{a}\mathsf{d}^{N}\}$. Then, the SDE describing the $N$-particle dynamics is such that, for $1\leq i\leq N$,
$X_{t}^{N,i}-X_{0}^{N,i}=\int_{0}^{t}\sigma(X_{u}^{N,i})dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}^{N}(X_{u}^{N,i},\sum_{j\neq i}^{N}\delta_{X_{u}^{N,j}})du.$ (2.55)
We impose reflecting boundary conditions on $\partial S_{R}$ in SDE (2.55) and hold the particles outside the region $\overline{S}_{R}=\{\lvert s\rvert\leq R\}$ fixed. Under these conditions, the SDE becomes
$\underline{X}_{R}^{N,i}(t)-\underline{X}_{R}^{N,i}(0)$ $=\int_{0}^{t}\sigma(\underline{X}_{R}^{N,i}(u))dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}^{N}(\underline{X}_{R}^{N,i}(u),\sum_{j\neq i}^{N}\delta_{\underline{X}_{R}^{N,j}(u)})du$
$\quad\quad\quad\quad+\int_{0}^{t}\frac{1}{2}\mathbf{n}_{R}\mathfrak{a}(\underline{X}_{R}^{N,i}(u))dL_{R}^{N,i}(u),$
$L_{R}^{N,i}(t)$ $=\int_{0}^{t}1_{\partial S_{R}}(\underline{X}_{R}^{N,i}(u))dL_{R}^{N,i}(u),\quad\text{for }i\leq\mathsf{s}(\overline{S}_{R}),$ (2.56)
$\underline{X}_{R}^{N,i}(t)-\underline{X}_{R}^{N,i}(0)$ $=0,\quad\text{for }\mathsf{s}(\overline{S}_{R})<i\leq N.$ (2.57)
@@
Here, $\mathbf{n}_{R}$ denotes the inward normal vector on the boundary $\partial S_{R}$, and $L_{R}^{N,i}$ is a non-negative, non-decreasing continuous process referred to as the local time on the boundary $\partial\overline{S}_{R}$ (cf. [7, pp. 251–256]).
We set the initial condition $\mathbf{X}_{R}^{N}(0)=\mathfrak{l}(\mathsf{s})$ for $\mu^{N}$-almost every $\mathsf{s}$, where $\mathfrak{l}=(\mathfrak{l}^{i})_{i}$ is the labeling introduced in (2.43). From (2.43) and (2.57), all particles outside the region $\overline{S}_{R}$ are held fixed.
Let $\{N_{\mathfrak{n}}\}_{n=1}^{\infty}$ be the sequence in ((A2)) such that $\mu^{N_{\mathfrak{n}}}$ converges to $\mu$. Taking the limit $\mathfrak{n}\to\infty$ in (2.56)–(2.57), we obtain the ISDE
$\underline{X}_{R}^{i}(t)-\underline{X}_{R}^{i}(0)$ $=\int_{0}^{t}\sigma(\underline{X}_{R}^{i}(u))dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}(\underline{X}_{R}^{i}(u),\sum_{j\neq i}^{\infty}\delta_{\underline{X}_{R}^{j}(u)})du$
$\quad\quad\quad\quad+\int_{0}^{t}\frac{1}{2}\mathbf{n}_{R}\mathfrak{a}(\underline{X}_{R}^{i}(u))dL_{R}^{i}(u),$
$L_{R}^{i}(t)$ $=\int_{0}^{t}1_{\partial S_{R}}(\underline{X}_{R}^{i}(u))dL_{R}^{i}(u),\quad\text{for }i\leq\mathsf{s}(\overline{S}_{R}),$ (2.58)
$\underline{X}_{R}^{i}(t)-\underline{X}_{R}^{i}(0)$ $=0,\quad\text{for }\mathsf{s}(\overline{S}_{R})<i<\infty.$ (2.59)
Remark 2.10.
i For $\mathbf{s}=(s^{i})_{i=1}^{N}$ such that $s^{i}\neq s^{j}$ for any $i\neq j$, SDE (2.56)–(2.57) with $\underline{\mathbf{X}}_{R}^{N}(0)=\mathbf{s}$ has a unique strong solution $\underline{\mathbf{X}}_{R}^{N}$. Indeed, the coefficients are smooth outside $\mathscr{N}=\cup_{i\neq j}^{N}\{s^{i}=s^{j}\}$ and $\underline{\mathbf{X}}_{R}^{N}$ does not hit $\mathscr{N}$ because $d\geq 2$. See Theorem 6.1.
ii The same as above holds for SDE (2.58)–(2.59) for $\mu$-a.s. $\mathsf{s}$. This claim follows from (2.33), (2.39), (2.40), and $\mathfrak{a}\in C_{b}^{2}(\mathbb{R}^{d})$. Hence, we denote by $\underline{\mathbf{X}}_{R}$ the unique strong solution for SDE (2.58)–(2.59).
We equip $C([0,\infty);\mathbb{R}^{d})$ with the topology of uniform convergence on each bounded interval and endow $C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$ with the corresponding product topology. With this topology, $C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$ becomes a Polish space.
Let $\mathfrak{l}_{N}=(\mathfrak{l}^{1},\dots,\mathfrak{l}^{N})$, where $\mathfrak{l}=(\mathfrak{l}^{i})_{i}$ is the labeling introduced in (2.43). Then
$\lim_{N\to\infty}\mathfrak{l}_{N}(\mathsf{s})=\mathfrak{l}(\mathsf{s})\quad\text{ for $\mu$-a.s.\,$\mathsf{s}$}.$
Let $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ be a solution of (2.56)–(2.57), and let $\underline{\mathbf{X}}_{R}$ and $\mathbf{X}$ be solutions of (2.58)–(2.59) and (2.41), respectively, as described in Remark 2.10 and Theorem 2.3. We impose
$\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}(0)=\mathfrak{l}_{N_{\mathfrak{n}}}(\mathsf{s}),\quad\underline{\mathbf{X}}_{R}(0)=\mathbf{X}(0)=\mathfrak{l}(\mathsf{s}).$ (2.60)
Therefore, using the label $\mathfrak{l}$, we couple the initial staring points of $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$, $\underline{\mathbf{X}}_{R}$, and $\mathbf{X}$.
Let $c_{2.4}$ be a positive constant. Let
@@
Let $\varphi d\mu$ be the random point field such that $(\varphi d\mu)(A)=\int_{A}\varphi d\mu$.
Theorem 2.5 (Finite-volume approximation).
Assume ((A1))((A3)). Let $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ be the solution of (2.56)–(2.57) with initial condition
$\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}(0)\sim_{\mathrm{law}}(\varphi d\mu)\circ(\mathfrak{l}_{N_{\mathfrak{n}}})^{-1}.$ (2.62)
Let $\underline{\mathbf{X}}_{R}$ and $\mathbf{X}$ be the solutions of (2.58)–(2.59) and (2.41) as in Remark 2.10 and Theorem 2.3 with initial condition
$\underline{\mathbf{X}}_{R}(0)=\mathbf{X}(0)\sim_{\mathrm{law}}(\varphi d\mu)\circ\mathfrak{l}^{-1}.$ (2.63)
Then, we have the following convergence in distribution in $C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$:
$\lim_{\mathfrak{n}\to\infty}\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ $=\underline{\mathbf{X}}_{R}\quad\text{ in distribution in }C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}},$ (2.64)
$\lim_{R\to\infty}\underline{\mathbf{X}}_{R}$ $=\mathbf{X}\quad\text{ in distribution in }C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}.$ (2.65)
Here, we regard $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ as $(\mathbb{R}^{d})^{\mathbb{N}}$-valued process by taking $\underline{X}_{R}^{N_{\mathfrak{n}},i}(t)=\underline{X}_{R}^{N_{\mathfrak{n}},i}(0)=\mathfrak{l}^{i}(\mathsf{s})$ for all $t\in[0,\infty)$ and $i>N_{\mathfrak{n}}$, where $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}=(\underline{X}_{R}^{N_{\mathfrak{n}},i})_{i\in\mathbb{N}}$.

### v4

@@ 2.2. Assumptions
((AC))ν $P\circ\mathsf{X}_{t}^{-1}\ll\nu$ for all $0<t<\infty$.
Condition ((SIN)) ensures that the dynamics remain in the admissible configuration space. Condition ((NBJ)) guarantees that, within any finite time interval, only finitely many particles enter a given bounded region. Condition ((AC))ν ensures absolute continuity of the time marginals with respect to the reference measure $\nu$.
In addition, we impose two further conditions. The condition ((MF)) is a mild measurability assumption, whereas ((IFC)) is a substantive condition whose formulation requires the finite-dimensional SDE scheme introduced later. They are introduced in Subsection 13.6 and Subsection 13.5, respectively. Precise definitions of strong solutions, unique strong solutions, and pathwise uniqueness are given in Definitions 13.1113.13.
The following Theorem 2.3 establishes the existence and uniqueness of a strong solution of (2.33) in the sense of Definition 13.13.
@@
ii For each $m\in\{0\}\cup\mathbb{N}$, the associated $m$-labeled process $\mathbf{X}^{[m]}$ is a $\mu_{\mathsf{a}}^{[m]}$-symmetric and conservative diffusion.
By Theorem 13.4, we obtain the following corollary of Theorem 2.3. Let $\xi_{\mathsf{a}}$ be an RPF such that $\xi_{\mathsf{a}}\ll\mu_{\mathsf{a}}$.
Corollary 2.1.
@@
From Theorems 2.12.3, we obtain the following corollary.
Corollary $2.$
Assume ((A1))((A3)). Then the conclusions of Theorem 2.2 and Theorem 2.3 apply to the ISDE (2.41):
$X_{t}^{i}-X_{0}^{i}=\int_{0}^{t}\sigma(X_{u}^{i})\,dB_{u}^{i}+\frac{\beta}{2}\int_{0}^{t}\mathrm{div\,}\mathfrak{a}(X_{u}^{i})\,du$
$+\frac{\beta}{2}\int_{0}^{t}\Big(-\nabla\Phi(X_{u}^{i})+\lim_{R\to\infty}\sum_{\begin{subarray}{c}\lvert X_{u}^{j}\rvert\leq R,\\ j\neq i\end{subarray}}\frac{X_{u}^{i}-X_{u}^{j}}{\lvert X_{u}^{i}-X_{u}^{j}\rvert^{d}}\Big)\mathfrak{a}(X_{u}^{i})\,du,\ i\in\mathbb{N}.$ (2.41)
@@
The analysis of fully labeled dynamics on the full label space $(\mathbb{R}^{d})^{\mathbb{N}}$ plays a crucial role in understanding infinite-particle systems. While the $m$-labeled dynamics $\mathbf{X}^{[m]}$ provide finite-dimensional projections that admit invariant measures, they capture only partial aspects of the underlying infinite system.
While the $m$-labeled dynamics $\mathbf{X}^{[m]}$ provide projections to the tame infinite-dimensional spaces that admit invariant measures, they capture only partial aspects of the underlying infinite system.
To obtain a complete description, it is necessary to go beyond finite labels and establish the existence of a genuine diffusion on $(\mathbb{R}^{d})^{\mathbb{N}}$. This step is indispensable, as it ensures that the dynamics are well defined for the entire infinite system rather than merely for restricted subsystems.
Historically, the construction of fully labeled dynamics has been regarded as a highly challenging problem. Even for determinantal processes such as the Dyson model, establishing such dynamics required delicate arguments based on random matrix theory. For general Coulomb systems, the situation is more severe: the long-range nature of the interaction and the absence of a classical Gibbsian structure prevent a direct application of Dirichlet form methods or invariant-measure techniques. Against this backdrop, proving the existence of fully labeled conservative diffusions represents a substantial advance.
In Theorem 2.4, we prove that the fully labeled process $\mathbf{X}$ constructed in Theorem 2.3 is an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued conservative diffusion. Unlike the $m$-labeled processes $\mathbf{X}^{[m]}$, the diffusion $\mathbf{X}$ is obtained without relying on the existence of an invariant measure, reflecting a fundamental difference between finite-label and full-label dynamics. Nevertheless, the result guarantees that the fully labeled dynamics form a legitimate stochastic process with infinite lifetime, thereby providing a rigorous foundation for the dynamical theory of Coulomb systems.
Let $\mathbf{S}_{\star\star}\subset(\mathbb{R}^{d})^{\mathbb{N}}$ be the subset defined in (11.14). Let $\mathbf{X}$ be the solution of the ISDE (2.33) constructed in Theorem 2.3. Let $\mathfrak{u}$ denote the unlabeled map as in (2.21).
@@
Assume ((A1))((A3)). Then $\mathbf{X}$ is a conservative diffusion with state space $\mathbf{S}_{\star\star}$ and $\mu(\mathfrak{u}(\mathbf{S}_{\star\star}))=1$.
@@
$P_{\mathbf{x}}:=P\bigl(\mathbf{X}\in\cdot\,\big|\,\mathbf{X}_{0}=\mathbf{x}\bigr)$
@@
admits a version such that the family $\{P_{\mathbf{x}}\}_{\mathbf{x}\in\mathbf{S}_{\star\star}}$ consists of probability measures on $C([0,\infty);\mathbf{S}_{\star\star})$ satisfying the strong Markov property and having infinite lifetime [5]. In particular, the diffusion law $P_{\mathbf{x}}$ is well defined for every $\mathbf{x}\in\mathbf{S}_{\star\star}$.
@@
The state space can, in principle, be extended to the whole $(\mathbb{R}^{d})^{\mathbb{N}}$ by freezing the dynamics outside $\mathbf{S}_{\star\star}$.
2.7. $N$-particle approximation: Theorem 2.5
We study finite-volume approximations of the infinite-particle dynamics. These approximations are realized in two complementary ways: by restricting the system to a bounded domain with reflecting boundary conditions (Proposition 12.1), and by truncating the system to $N$ particles (Theorem 2.5). These theorems rigorously establish the convergence of these finite-volume approximations to the infinite-dimensional dynamics.
Let $\mathsf{d}^{N}$ be the logarithmic derivative of $\mu^{N}$. By (1.3) and ((A1)),
$\mathsf{d}^{N}({x},\mathsf{s})=-\beta\Big(\nabla\Phi^{N}({x})+\sum_{i=1}^{N-1}\nabla_{{x}}\Psi^{N}({x},s^{i})\Big),\quad{x}\in O_{N}.$
@@
Set $\mathfrak{b}^{N}=\frac{1}{2}(\mathrm{div\,}\mathfrak{a}+\mathsf{d}^{N}\mathfrak{a})$. Then the SDE describing the $N$-particle dynamics $\mathbf{X}^{N}=(X^{N,i})_{i=1}^{N}$ is given by, for $1\leq i\leq N$,
$X_{t}^{N,i}-X_{0}^{N,i}$ $=\int_{0}^{t}\sigma(X_{u}^{N,i})\,dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}^{N}\Big(X_{u}^{N,i},\sum_{j\neq i}^{N}\delta_{X_{u}^{N,j}}\Big)du$
$\quad\quad\quad+\int_{0}^{t}\frac{1}{2}\mathbf{n}_{N}\mathfrak{a}(X_{u}^{N,i})\,dL_{N}^{N,i}(u),$
$L_{N}^{N,i}(t)$ $=\int_{0}^{t}1_{\partial O_{N}}(X_{u}^{N,i})\,dL_{N}^{N,i}(u).$ (2.42)
@@
Here $\frac{1}{2}\mathbf{n}_{N}\mathfrak{a}$ denotes the inward unit normal vector on $\partial O_{N}$, and $L_{N}^{N,i}$ is a nonnegative, nondecreasing continuous process, called the local time on the boundary $\partial O_{N}$ (cf. [8, pp. 251–256]). If $O_{N}=\mathbb{R}^{d}$, no local time term appears in (2.42). The SDE then becomes
$X_{t}^{N,i}-X_{0}^{N,i}=\int_{0}^{t}\sigma(X_{u}^{N,i})\,dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}^{N}\Big(X_{u}^{N,i},\sum_{j\neq i}^{N}\delta_{X_{u}^{N,j}}\Big)du.$
We impose
$\mathbf{X}^{N}_{0}=\mathfrak{l}_{N}(\mathsf{s}),\qquad\mathbf{X}_{0}=\mathfrak{l}(\mathsf{s}).$ (2.43)
Here $\mathfrak{l}=(\mathfrak{l}^{i})_{i}$ denotes the labeling map introduced in (2.22), and $\mathfrak{l}_{N}=(\mathfrak{l}^{i})_{i=1}^{N}$. The initial conditions of these solutions are coupled through the labeling map.
Let $c_{2.5}$ be a positive constant. Let
@@
Define the RPF $\varphi d\mu$ by $(\varphi d\mu)(A)=\int_{A}\varphi d\mu$ for $A\in\mathscr{B}(\mathsf{S})$. Let $\mathbf{X}^{N}$ and $\mathbf{X}$ be the (unique strong) solutions of SDEs (2.42) and (2.33), respectively. We assume:
$\mathbf{X}^{N}_{0}\mathop{\sim}\limits_{\mathrm{law}}(\varphi d\mu)\circ(\mathfrak{l}_{N})^{-1},\quad\mathbf{X}_{0}\mathop{\sim}\limits_{\mathrm{law}}(\varphi d\mu)\circ\mathfrak{l}^{-1}.$ (2.45)
By setting $X_{t}^{N,i}=\mathfrak{l}^{i}(\mathsf{s})$ for all $t\geq 0$ and $i>N$, we regard $\mathbf{X}^{N}=(X^{N,i})$ as a $W(\mathbb{R}^{d})^{\mathbb{N}}$-valued random variable.
Theorem 2.5 (NN-particle approximation).
Assume ((A1))((A3)) and (2.44)–(2.45). Then
$\lim_{\mathfrak{n}\to\infty}\mathbf{X}^{N_{\mathfrak{n}}}$ $=\mathbf{X}\quad\text{ in law in }W(\mathbb{R}^{d})^{\mathbb{N}}.$